# Learning Mean Field Control on Sparse Graphs

### v1

@@ 1 Introduction
Despite the rapid developments in the field of multi-agent reinforcement learning (MARL) over the last years, systems with many agents remain hard to solve in general (Canese et al., 2021; Gronauer & Diepold, 2022). Mean field games (MFGs) (Caines et al., 2006; Lasry & Lions, 2007) and mean field control (MFC) (Andersson & Djehiche, 2011; Bensoussan et al., 2013) are a promising way to model large agent problems in a computationally tractable manner and to provide a solid theoretical framework at the same time. While MFGs consider competitive agent populations, the focus of MFC are cooperative scenarios where agents optimize a common goal. The idea of MFC and MFGs is to abstract large, homogeneous crowds of small agents into a single probability distribution, the mean field (MF). While MFC and MFGs have been used in various areas ranging from pedestrian flows (Bagagiolo et al., 2019; Achdou & Laurière, 2020) to finance (Carmona & Delarue, 2018; Carmona & Laurière, 2023) and oil production (Bauso et al., 2016), the assumption of indistinguishable agents is not fulfilled in many applications.
@@
A particularly important class of MARL problems are those with many connected agents. Initially, these agent networks were modeled by combining the graph theoretical concept of graphons (Lovász, 2012) with MFGs, resulting in graphon MFGs (GMFGs) (Caines & Huang, 2019, 2021; Cui & Koeppl, 2022; Zhang et al., 2024) and graphon MFC (Hu et al., 2023b). Since GMFGs only model often unrealistic dense graphs, subsequently mean field models based on Lp graphons (Borgs et al., 2018b, 2019) and graphexes (Veitch & Roy, 2015; Caron & Fox, 2017; Borgs et al., 2018a) were developed, called LPGMFGs and GXMFGs, respectively (Fabian et al., 2023, 2024). While these models facilitate learning algorithms in moderately sparse networks, they exclude sparser topologies. Formally, (LP)GMFGs and GXMFGs are designed exclusively for graphs with expected average degree going to infinity which, for example, excludes power laws with a coefficient above two.
@@
The learning literature contains various approaches to finding optimal behavior in MFGs and MFC, see Laurière et al. (2022a) for an overview. For example, Subramanian et al. (2022) develop a decentralized learning algorithm for MFGs where agents are able to independently learn policies, while Guo et al. (2019, 2023) focus on Q-learning methods for general MFGs. Various MFC learning approaches exist (Ruthotto et al., 2020; Carmona et al., 2023; Gu et al., 2023), but we are aware of only one work by Hu et al. (2023b) which learns policies for MFC on dense networks, but not on sparse ones.
@@
Many empirical networks of high practical relevance are considerably sparser than the topologies covered by (LP)GMFGs and GXMFGs. Examples of sparse empirical networks which at least to some extent follow power laws with coefficients between two and three include the internet (Vázquez et al., 2002), coauthorship graphs (Goh et al., 2002) and biological networks (Dorogovtsev & Mendes, 2002). These topologies are particularly challenging to analyze (Dorogovtsev et al., 2002), especially in combination with particle dynamics (Lacker et al., 2023). Most important in our context, the high fraction of low degree nodes in these networks renders classical mean field approximations and their LPGMFG and GXMFG extensions highly inaccurate.
@@
Both LPGMFGs and GXMFGs assume that the average expected degree diverges to infinity to ensure that the neighborhoods of crucial agents are accurately approximated by a mean field. Since this assumption is inaccurate for various empirical networks, we require a different modeling and learning approach. To learn policies for such very sparse networks, we employ a suitable graph theoretical convergence principle, local weak convergence (Van Der Hofstad, 2024). Many established graph theoretical models like the configuration model (Bollobás, 1980), the Barabási-Albert model (Barabási & Albert, 1999) and the Chung-Lu model (Chung & Lu, 2006) can model graph sequences converging in the local weak sense. Most importantly in our context, this includes graph sequences with finite average expected degree and power law networks with a coefficient above two, which are neither covered by Lp graphons nor graphexes.
@@ 2 Locally Weak Converging Graphs
In the following, let $(G_{N})_{N\in\mathbb{N}}=(V_{N},E_{N})_{N\in\mathbb{N}}$ be a growing sequence of random graphs where $V_{N}$ denotes the vertex set and $E_{N}$ is the edge set of the corresponding graph $G_{N}$. In this paper, we focus on growing graph sequences where the expected average degree remains finite in the limit while the degree variance may diverge to infinity. To formalize the properties of the graph sequences we are focusing on, we first require a suitable graph convergence concept. We choose local weak convergence in probability which means that local node neighborhoods converge to neighborhoods in a limiting model. The definition below states local weak convergence, for details see e.g. Lacker et al. (2023).
@@
A sequence of finite graphs $(G_{N})_{N}$ converges in probability in the local weak sense to $G\in\mathcal{G}^{*}$ if for all continuous and bounded functions $f:\mathcal{G}^{*}\to\mathbb{R}$
$\lim_{N\to\infty}\frac{1}{N}\sum_{i\in[N]}f(C_{v_{i}}(G_{N}))=\mathbb{E}[f(G)]\quad\textrm{in probability},$
@@
The sequence $(G_{N})_{N\in\mathbb{N}}$ converges in probability in the local weak sense to some $G\in\mathcal{G}^{*}$.
The class of random graph sequences fulfilling Assumption 2.2 covers many famous graph theoretical frameworks which are frequently used in the literature. We will briefly discuss three particularly important types of these models, namely configuration models, preferential attachment models and Chung-Lu graphs. We point to Van Der Hofstad (2024) for an extensive introduction and theoretical details and for more random graph models converging in the local weak sense.
@@
The configuration model (Bender & Canfield, 1978; Bollobás, 1980; Molloy & Reed, 1995, 1998) (CM) is arguably one of the most established random graph models. The basic mechanism of the CM is to start with a fixed and arbitrary degree sequence. Then, a multigraph is randomly generated with the prescribed degree distribution which means that the graph can contain self-edges and double edges between pairs of nodes.
The CM is known to converge under suitable and moderate assumptions in the local weak sense in probability (Van Der Hofstad, 2024, Theorem 4.1). However, the CM generates multigraphs instead of simple graphs and the number of multiedges increases drastically as the vertex degrees increase (Bollobás, 1998). Consequently, the CM is suboptimal for generating simple graphs with a significant fraction of high degree nodes such as power law networks.
@@
To model random graphs with power law features, Barabási & Albert (1999) introduced the famous Barabási-Albert (BA) model. The original BA model was subsequently extended in various ways for different applications, see Piva et al. (2021) for an overview. The BA model depicts graphs that grow over time by adding new nodes and edges to the topology. Since new nodes are more likely to be connected to highly connected nodes in the current graph, these models are often referred to as preferential attachment models.
Preferential attachment models under suitable conditions converge in the local weak sense in probability, see Van Der Hofstad (2024, Theorem 5.8) for details. The BA model, for example, generates power law networks with a coefficient of exactly three (Bollobás et al., 2001). In many applications, however, it is beneficial to also consider graphs with power law coefficient deviating from three to capture different empirical graph topologies.
@@
The Chung-Lu (CL) random graph model (Aiello et al., 2000, 2001; Chung & Lu, 2002, 2006) provides an efficient way to model large, sparse networks (Fasino et al., 2021). To generate a random CL graph with $N\in\mathbb{N}$ nodes, first specify a weight vector $\boldsymbol{w}\in\mathbb{R}_{+}^{N}$ with one weight $w_{i}\in\mathbb{R}_{+}$ for each node $i\in\{1,\ldots,N\}$. Then, two nodes $i$ and $j$ are connected with probability $w_{i}\cdot w_{j}/\bar{w}$, independently of all other node pairs and with normalization factor $\bar{w}\coloneqq\sum_{1\leq k\leq N}w_{k}$. Intuitively, a node with high weight is more likely to have many connections than a node with small weight.
Sequences of CL graphs fulfill Assumption 2.2 under mild technical assumptions, see Van Der Hofstad (2024, Theorem 3.18) for a formal statement. Most importantly, the average expected degree has to converge to a finite limit, which perfectly aligns with our goal to model very sparse networks. Throughout the paper, we use the running example of power law degree distributions with coefficient $\gamma>2$ observed in many real world networks to some extent (Newman, 2003; Kaufmann & Zweig, 2009; Newman et al., 2011).
@@
Large, sparse power law networks of the above form can be efficiently generated by the CL framework (Fasino et al., 2021). Note that our methods apply to all distributions meeting Assumption 2.2 and perform well on many empirical networks, as shown in the next sections.
@@
Local weak converging graph sequences such as those generated by CM, BM or CL can model sparser, and thus often more realistic topologies than those captured by Lp graphons and graphexes. Figure 1 provides an illustration of how Lp graphons, graphexes, and CL graphs compare to a real world subsampled YouTube network (Mislove, 2009; Kunegis, 2013). Mathematically, both Lp graphons and graphexes are limited to graph sequences were the average degree diverges to infinity. Locally weak converging graph sequences, on the other hand, can capture sparse and often more realistic topologies. The usefulness of models like the CM, BA model and CL graphs is reflected in their frequent use in various research areas. However, formulating a mean field approach based on these graph theoretical models is challenging due to their high number of low degree nodes.
Figure 1: Four networks, first two generated by an Lp graphon and graphex, third is a CL graph and fourth is a real subsampled YouTube (YT) network (Mislove, 2009; Kunegis, 2013), highly connected nodes are depicted larger. Each network has around 14.5k nodes and 13k edges, except graphex has around 16.5k edges; all networks are plotted in the prefuse force directed layout (software: cytoscape). While the Lp graphon graph lacks sufficiently many high degree nodes, the tail of the graphex degree distribution is too heavy. In contrast, the CL graph is qualitatively close to the real YT network.
@@ 3 The Finite Model and Its Limit
Denote by $\mathcal{P}(\mathcal{X})$ the set of probability distributions over a finite set $\mathcal{X}$ and define $[N]\coloneqq\{1,\ldots,N\}$ for any $N\in\mathbb{N}$.
@@
$\mu_{t}^{N,k}\coloneqq\frac{1}{|V_{N}^{k}|}\sum_{i\in[N]:v_{i}\in V_{N}^{k}}\delta_{X_{i,t}^{N}}\in\mathcal{P}(\mathcal{X}),$
@@ 4 The Two Systems Approximation
In limiting systems on sparse graphs, the state evolution and optimal policy of an agent potentially depend on the entire network (Lacker & Soret, 2022). Calculating $P_{\boldsymbol{\pi}}\left(\mathbb{G}_{t}^{k}\left(\boldsymbol{\mu}^{\prime}_{t}\right)=G\mid x_{t}=x\right)$ at time $t\in\mathcal{T}$ in the limiting system requires all possible $t$-hop neighborhood degree-state distributions where $t$-hop neighborhoods include all agents with a distance of at most $t$ edges to the initial agent. Unfortunately, by Lemma 4.1 the number of $t$-hop neighborhoods grows at least exponentially with the degree $k$ in important classes of locally weak converging graphs sequences, such as CL graphs with power laws above two.
@@
Heuristic 1 is a good approximation for some sequences of locally weak converging graphs, such as CL graphs (Jackson et al., 2008, Chapter 4), and thus reasonable in our setup. The idea of Heuristic 1 is the following: if one fixes any node $v^{\prime}\in V$ and considers its neighbors, high degree nodes are more likely to be connected to $v^{\prime}$ than lowly connected ones. Instead of the overall degree distribution, we thus weight each probability by its degree and normalize accordingly. The result is an approximate neighbor degree distribution accounting for the increased probability of highly connected neighbors.
@@
$\hat{\mathbb{G}}_{t}^{\infty}(\hat{\boldsymbol{\mu}})$ $\coloneqq\frac{1}{\mathbb{E}[\deg(v)]}\left(\sum_{k=k^{*}+1}^{\infty}kP(\deg(v)=k)\right)\hat{\mu}_{t}^{\infty}$
@@
$\,\sum_{x,u}\hat{\mu}^{\infty}_{t}(x)\pi_{t}^{\infty}\left(u\mid x,\hat{\mathbb{G}}_{t}^{\infty}(\boldsymbol{\mu}^{\prime})\right)P\left(\cdot\,\mid x,u,\hat{\mathbb{G}}_{t}^{\infty}(\boldsymbol{\mu}^{\prime})\right),$
where all agents with more than $k^{*}$ connections follow the same policy $\pi_{t}^{\infty}\in\mathcal{P}(\mathcal{U})^{\mathcal{T}\times\mathcal{X}\times\mathcal{P}(\mathcal{X})}$ and where the sum is over all $(x,u)\in\mathcal{X}\times\mathcal{U}$. The approximate neighborhood of an agent with degree $k\in[k^{*}]$ at each time $t\in\mathcal{T}$ is sampled from $\hat{\mathbb{G}}_{t}^{k}(\hat{\boldsymbol{\mu}})\sim\mathrm{Mult}(k,\hat{\mathbb{G}}_{t}^{\infty}(\hat{\boldsymbol{\mu}}))$, i.e. $\hat{\mathbb{G}}_{t}^{k}(\hat{\boldsymbol{\mu}})$ is multinomial with $k$ trials and probabilities $\hat{\mathbb{G}}_{t}^{\infty}(\hat{\boldsymbol{\mu}})(x)$ for each $x\in\mathcal{X}$. Using Heuristic 1, the approximation yields for each $k\in[k^{*}]$ the MF forward dynamics
@@
of the finite agent neighborhoods in the LWMFC model where we use the abbreviated notation $p_{j,m}(\boldsymbol{a}_{2})\coloneqq\left(P\left(\deg(v)=m\mid(v^{\prime},v)\in E\right)\mu_{t}^{m}(s_{j})\right)^{a_{jm}}$. Here, the idea is to go beyond the previous multinomial assumption $\hat{\mathbb{G}}_{t}^{k}(\hat{\boldsymbol{\mu}})\sim\mathrm{Mult}(k,\hat{\mathbb{G}}_{t}^{\infty}(\hat{\boldsymbol{\mu}}))$ and to use state-degree neighborhood distributions $\boldsymbol{a}_{2}\in\boldsymbol{\mathcal{A}}^{k}_{2}(G^{\prime},c)$ and state-state-degree neighborhood distributions $\boldsymbol{a}_{3}\in\boldsymbol{\mathcal{A}}^{k}_{3}(G,G^{\prime},c)$ to capture agents changing from $x\in\mathcal{X}$ to $x^{\prime}\in\mathcal{X}$ at a time step. We provide the extensive approximation derivation and corresponding definitions of sets like $\boldsymbol{\mathcal{A}}^{k}_{2}(G^{\prime},c)$ and $\boldsymbol{\mathcal{A}}^{k}_{3}(G,G^{\prime},c)$ in Appendix B. As we will see in the following, the extensive approximation often shows a moderately higher accuracy than our first approximation. However, the accuracy boost entails a significantly higher computational complexity due to multiple sums over sets like $\boldsymbol{\mathcal{A}}^{k}_{2}(G^{\prime},c)$ and $\boldsymbol{\mathcal{A}}^{k}_{3}(G,G^{\prime},c)$. Thus, our first approximation is more practical since it combines reasonable accuracy with low computational complexity while the extensive approximation is computationally too expensive for our purposes.

### v2

@@ 1 Introduction
Despite the rapid developments in the field of multi-agent reinforcement learning (MARL) over the last years, systems with many agents remain hard to solve in general (canese2021multi; gronauer2022multi). Mean field games (MFGs) (caines2006large; lasry2007mean) and mean field control (MFC) (andersson2011maximum; bensoussan2013mean) are a promising way to model large agent problems in a computationally tractable manner and to provide a solid theoretical framework at the same time. While MFGs consider competitive agent populations, the focus of MFC are cooperative scenarios where agents optimize a common goal. The idea of MFC and MFGs is to abstract large, homogeneous crowds of small agents into a single probability distribution, the mean field (MF). While MFC and MFGs have been used in various areas ranging from pedestrian flows (bagagiolo2019optimal; Achdou2020) to finance (carmona2018probabilistic; carmona2023deep) and oil production (bauso2016robust), the assumption of indistinguishable agents is not fulfilled in many applications.
@@
A particularly important class of MARL problems are those with many connected agents. Initially, these agent networks were modeled by combining the graph theoretical concept of graphons (lovasz2012large) with MFGs, resulting in graphon MFGs (GMFGs) (caines2019graphon; caines2021graphon; cuilearning; zhang2024learning) and graphon MFC (hu2023graphon). Since GMFGs only model often unrealistic dense graphs, subsequently mean field models based on Lp graphons (borgs2018lp; borgs2019L) and graphexes (veitch2015class; caron2017sparse; borgs2018sparse) were developed, called LPGMFGs and GXMFGs, respectively (fabian2023learning; fabian2024learning). While these models facilitate learning algorithms in moderately sparse networks, they exclude sparser topologies. Formally, (LP)GMFGs and GXMFGs are designed exclusively for graphs with expected average degree going to infinity which, for example, excludes power laws with a coefficient above two.
@@
The learning literature contains various approaches to finding optimal behavior in MFGs and MFC, see lauriere2022learning for an overview. For example, subramanian2022decentralized develop a decentralized learning algorithm for MFGs where agents independently learn policies, while guo2019learning; guo2023general focus on Q-learning methods for general MFGs. Various MFC learning approaches exist (ruthotto2020machine; carmona2023model; gu2023dynamic), but we are aware of only one work by hu2023graphon which learns policies for MFC on dense networks, but not on sparse ones.
@@
Many empirical networks of high practical relevance are considerably sparser than the topologies covered by (LP)GMFGs and GXMFGs. Examples of sparse empirical networks which at least to some extent follow power laws with coefficients between two and three include the internet (vazquez2002large), coauthorship graphs (goh2002classification) and biological networks (dorogovtsev2002evolution). These topologies are particularly challenging to analyze (dorogovtsev2002pseudofractal), especially in combination with particle dynamics (lacker2023local). Most important in our context, the high fraction of low degree nodes in these networks renders classical mean field approximations and their LPGMFG and GXMFG extensions highly inaccurate.
@@
Both LPGMFGs and GXMFGs assume that the average expected degree diverges to infinity to ensure that the neighborhoods of crucial agents are accurately approximated by a mean field. Since this assumption is inaccurate for various empirical networks, we require a different modeling and learning approach. To learn policies for such very sparse networks, we employ a suitable graph theoretical convergence principle, local weak convergence (van2024random). Many established graph theoretical models like the configuration model (bollobas1980probabilistic), the Barabási-Albert model (barabasi1999emergence) and the Chung-Lu model (chung2006complex) can model graph sequences converging in the local weak sense. Most importantly in our context, this includes graph sequences with finite average expected degree and power law networks with a coefficient above two, which are neither covered by Lp graphons nor graphexes.
@@ 2 Locally Weak Converging Graphs
In the following, let $(G_{N})_{N\in\mathbb{N}}=(V_{N},E_{N})_{N\in\mathbb{N}}$ be a growing sequence of random graphs where $V_{N}$ denotes the vertex set and $E_{N}$ is the edge set of the corresponding graph $G_{N}$. In this paper, we focus on growing graph sequences where the expected average degree remains finite in the limit while the degree variance may diverge to infinity. To formalize the properties of the graph sequences we are focusing on, we first require a suitable graph convergence concept. We choose local weak convergence in probability which means that local node neighborhoods converge to neighborhoods in a limiting model. The definition below states local weak convergence, for details see e.g. lacker2023local.
@@
A sequence of finite graphs $(G_{N})_{N}$ converges in probability in the local weak sense to a random element $G$ of $\mathcal{G}^{*}$ if for all continuous and bounded functions $f:\mathcal{G}^{*}\to\mathbb{R}$
$\lim_{N\to\infty}\frac{1}{N}\sum_{i\in[N]}f(C_{v_{i}}(G_{N}))=\mathbb{E}[f(G)]\quad\textrm{in probability}$
@@
The sequence $(G_{N})_{N\in\mathbb{N}}$ converges in probability in the local weak sense to some random element $G$ of $\mathcal{G}^{*}$.
The class of random graph sequences fulfilling Assumption 2.2 covers many famous graph theoretical frameworks which are frequently used in the literature. We will briefly discuss three particularly important types of these models, namely configuration models, preferential attachment models and Chung-Lu graphs. We point to van2024random for an extensive introduction and theoretical details and for more random graph models converging in the local weak sense.
@@
The configuration model (bender1978asymptotic; bollobas1980probabilistic; molloy1995critical; molloy1998size) (CM) is arguably one of the most established random graph models. The basic mechanism of the CM is to start with a fixed and arbitrary degree sequence. Then, a multigraph is randomly generated with the prescribed degree distribution which means that the graph can contain self-edges and double edges between pairs of nodes.
The CM is known to converge under suitable and moderate assumptions in the local weak sense in probability (van2024random, Theorem 4.1). However, the CM generates multigraphs instead of simple graphs and the number of multiedges increases drastically as the vertex degrees increase (bollobas1998random). Consequently, the CM is suboptimal for generating simple graphs with a significant fraction of high degree nodes such as power law networks.
@@
To model random graphs with power law features, barabasi1999emergence introduced the famous Barabási-Albert (BA) model. The original BA model was subsequently extended in various ways for different applications, see piva2021networks for an overview. The BA model depicts graphs that grow over time by adding new nodes and edges to the topology. Since new nodes are more likely to be connected to highly connected nodes in the current graph, these models are often referred to as preferential attachment models.
Preferential attachment models under suitable conditions converge in the local weak sense in probability, see van2024random for details. The BA model, for example, generates power law networks with a coefficient of exactly three (bollobas2001degree). In many applications, however, it is beneficial to also consider graphs with power law coefficient deviating from three to capture different empirical graph topologies.
@@
The Chung-Lu (CL) random graph model (aiello2000random; aiello2001random; chung2002connected; chung2006complex) provides an efficient way to model large, sparse networks (fasino2021generating). To generate a random CL graph with $N\in\mathbb{N}$ nodes, first specify a weight vector $\boldsymbol{w}\in\mathbb{R}_{+}^{N}$ with one weight $w_{i}\in\mathbb{R}_{+}$ for each node $i\in\{1,\ldots,N\}$. Then, two nodes $i$ and $j$ are connected with probability $w_{i}\cdot w_{j}/\bar{w}$, independently of all other node pairs and with normalization factor $\bar{w}\coloneqq\sum_{1\leq k\leq N}w_{k}$. Intuitively, a node with high weight is more likely to have many connections than a node with small weight.
Sequences of CL graphs fulfill Assumption 2.2 under mild technical assumptions, see van2024random for a formal statement. Most importantly, the average expected degree has to converge to a finite limit, which perfectly aligns with our goal to model very sparse networks. Throughout the paper, we use the running example of power law degree distributions with coefficient $\gamma>2$ observed in many real world networks to some extent (newman2003structure; kaufmann2009modeling; newman2011structure).
@@
Large, sparse power law networks of the above form can be efficiently generated by the CL framework (fasino2021generating). Note that our methods apply to all distributions meeting Assumption 2.2 and perform well on many empirical networks, as shown in the next sections.
@@
Local weak converging graph sequences such as those generated by CM, BM or CL can model sparser, and thus often more realistic topologies than those captured by Lp graphons and graphexes. Figure 1 provides an illustration of how Lp graphons, graphexes, and CL graphs compare to a real world subsampled YouTube network (mislove2009online; kunegis2013konect). Mathematically, both Lp graphons and graphexes are limited to graph sequences were the average degree diverges to infinity. Locally weak converging graph sequences, on the other hand, can capture sparse and often more realistic topologies. The usefulness of models like the CM, BA model and CL graphs is reflected in their frequent use in various research areas. However, formulating a mean field approach based on these graph theoretical models is challenging due to their high number of low degree nodes.
Figure 1: Four networks, first two generated by an Lp graphon and graphex, third is a CL graph and fourth is a real subsampled YouTube (YT) network (mislove2009online; kunegis2013konect), highly connected nodes are depicted larger. Each network has around 14.5k nodes and 13k edges, except graphex has around 16.5k edges; all networks are plotted in the prefuse force directed layout (software: cytoscape). While the Lp graphon graph lacks sufficiently many high degree nodes, the tail of the graphex degree distribution is too heavy. In contrast, the CL graph is qualitatively close to the real YT network.
@@ 3 The Finite Model and Its Limit
Denote by $\mathcal{P}(\mathcal{X})$ the set of probability distributions over a finite set $\mathcal{X}$ and define $N]\coloneqq\{1,\ldots,N\}$ for any $N\in\mathbb{N}$.
@@
$\mu_{t}^{N,k}\coloneqq\frac{1}{|V_{N}^{k}|}\sum_{i\in[N]:v_{i}\in V_{N}^{k}}\delta_{X_{i,t}^{N}}\in\mathcal{P}(\mathcal{X})$
@@ 4 The Two Systems Approximation
In limiting systems on sparse graphs, the state evolution and optimal policy of an agent potentially depend on the entire network (lacker2022case). Calculating $P_{\boldsymbol{\pi}}\left(\mathbb{G}_{t}^{k}\left(\boldsymbol{\mu}^{\prime}_{t}\right)=G\mid x_{t}=x\right)$ at time $t\in\mathcal{T}$ in the limiting system requires all possible $t$-hop neighborhood degree-state distributions where $t$-hop neighborhoods include all agents with a distance of at most $t$ edges to the initial agent. Unfortunately, by Lemma 4.1 the number of $t$-hop neighborhoods grows at least exponentially with the degree $k$ in important classes of locally weak converging graphs sequences, such as CL graphs with power laws above two.
@@
Heuristic 1 is a good approximation for some sequences of locally weak converging graphs, such as CL graphs (jackson2008social, Chapter 4), and thus reasonable in our setup. The idea of Heuristic 1 is the following: if one fixes any node $v^{\prime}\in V$ and considers its neighbors, high degree nodes are more likely to be connected to $v^{\prime}$ than lowly connected ones. Instead of the overall degree distribution, we thus weight each probability by its degree and normalize accordingly. The result is an approximate neighbor degree distribution accounting for the increased probability of highly connected neighbors.
@@
[…]
@@
$\,\sum_{x,u}\hat{\mu}^{\infty}_{t}(x)\pi_{t}^{\infty}\left(u\mid x,\hat{\mathbb{G}}_{t}^{\infty}(\boldsymbol{\mu}^{\prime})\right)P\left(\cdot\,\mid x,u,\hat{\mathbb{G}}_{t}^{\infty}(\boldsymbol{\mu}^{\prime})\right)$
[…]
@@
of the finite agent neighborhoods in the LWMFC model where we use the abbreviated notation $p_{j,m}(\boldsymbol{a}_{2})\coloneqq\left(P\left(\deg(v)=m\mid(v^{\prime},v)\in E\right)\mu_{t}^{m}(s_{j})\right)^{a_{jm}}$. Here, the idea is to go beyond the previous multinomial assumption $)\hat{\mathbb{G}}_{t}^{k}(\hat{\boldsymbol{\mu}})\sim\mathrm{Mult}(k,\hat{\mathbb{G}}_{t}^{\infty}(\hat{\boldsymbol{\mu}}$)) and to use state-degree neighborhood distributions $\boldsymbol{a}_{2}\in\boldsymbol{\mathcal{A}}^{k}_{2}(G^{\prime},c)$ and state-state-degree neighborhood distributions $\boldsymbol{a}_{3}\in\boldsymbol{\mathcal{A}}^{k}_{3}(G,G^{\prime},c)$ to capture agents changing from $x\in\mathcal{X}$ to $x^{\prime}\in\mathcal{X}$ at a time step. We provide the extensive approximation derivation and corresponding definitions of sets like $\boldsymbol{\mathcal{A}}^{k}_{2}(G^{\prime},c)$ and $\boldsymbol{\mathcal{A}}^{k}_{3}(G,G^{\prime},c)$ in Appendix B. As we will see in the following, the extensive approximation often shows a moderately higher accuracy than our first approximation. However, the accuracy boost entails a significantly higher computational complexity due to multiple sums over sets like $\boldsymbol{\mathcal{A}}^{k}_{2}(G^{\prime},c)$ and $\boldsymbol{\mathcal{A}}^{k}_{3}(G,G^{\prime},c)$. Thus, our first approximation is more practical since it combines reasonable accuracy with low computational complexity while the extensive approximation is computationally too expensive for our purposes.